# Selecting informative conformal prediction sets with false coverage rate control

### v1

@@ 2.1 Conformal prediction sets
where $\mathbf{p}=(p^{(y)}_{i},y\in\mathcal{Y},i\in[m]\}$ is a collection of conformal $p$-values satisfying the following super-uniform guarantee:
@@
where the above probability is compute either in the iid ($Y_{n+i}$ random) or conditional ($Y_{n+i}$ fixed) model. Super-uniformity (5) implies directly that $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ in (4) provides $1-\alpha$ coverage for $Y_{n+i}$, that is,
@@
The $p$-value family is built from the calibration and test sample by using non-conformity score functions $S_{y}:x\in\mathbb{R}^{d}\mapsto\mathbb{R}$, $y\in\mathcal{Y}$, measuring the inadequacy between $y$ and the prediction at point $x$. Importantly, we follow a split/inductive conformal approach, where the score functions have been computed from an independent training data sample so that they can be considered as fixed here [^4] Note that we can relax slightly this assumption: our theory also allows this function to depend on the calibration plus test samples in an exchangeable way, see Assumption 6 in the SM. (and all probabilities/expectations are taken conditional on that training sample).
@@
The score functions $S_{y}(\cdot)$, $y\in\mathcal{Y}$, have been computed from an independent training data sample and the computed scores $(S_{Y_{i}}(X_{i}),i\in[n+m])$ have no ties almost surely. Moreover, when $\mathcal{Y}=\mathbb{R}$, the score function is regular in the following sense: for every $x\in\mathbb{R}^{d}$, the function $y\in\mathbb{R}\mapsto S_{y}(x)\in\mathbb{R}$ is right continuous with left limits.
@@
An equivalent description of the prediction set $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ in (4) can be given as a score level set, with a threshold depending on the calibration sores. Formally, we have
@@
[…]
@@ 2.2 $\mathcal{I}$-adjusted $p$-values
    If a prediction set is informative, then all the prediction sets it contains are also informative, that is,
    • Classification: for all $\mathcal{C},\mathcal{C}^{\prime}\subset[K]$, if $\mathcal{C}^{\prime}\subset\mathcal{C}$ and $\mathcal{C}\in\mathcal{I}$, then $\mathcal{C}^{\prime}\in\mathcal{I}$.
    • Regression: for all $\mathcal{C},\mathcal{C}^{\prime}$ intervals of $\mathbb{R}$, if $\mathcal{C}^{\prime}\subset\mathcal{C}$ and $\mathcal{C}\in\mathcal{I}$, then $\mathcal{C}^{\prime}\in\mathcal{I}$.
@@
    (Only in the regression case) For all $\alpha\in(0,1)$, almost surely, $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ is an interval of $\mathbb{R}$.
@@
with by convention $q_{i}=1$ if the set is empty. Assumption 2 can be easily checked for Examples 1.1 and 1.2, with an explicit expression for $q_{i}$’s.
@@
For regression, that is $\mathcal{Y}=\mathbb{R}$ (see § 4 for more details):
  1. 1.
    Intervals excluding $\mathcal{Y}_{0}=[a,b]$: $q_{i}=\max_{y\in[a,b]}p^{(y)}_{i}$ (if this maximum exists).
  2. 2.
    Length-restricted intervals $I$ (with $|I|\leq 2\lambda_{0}$):
    • for the locally weighted score function $S_{y}(x)=|y-\mu(x)|/\sigma(x)$,
      $q_{i}=(n+1)^{-1}\Big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})>\lambda_{0}/\sigma(X_{n+i})\}}\Big{)};$ (12)
    • for the quantile-based score function $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$,
      $q_{i}=(n+1)^{-1}\Big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})>\lambda_{0}-(q_{\beta_{1}}(X_{n+i})-q_{\beta_{0}}(X_{n+i}))/2\}}\Big{)}.$ (13)
@@
For classification, that is, $\mathcal{Y}=[K]$:
  1. 1.
    Excluding one-class: $q_{i}=p^{(y_{0})}_{i}$; or excluding several classes: $q_{i}=\max_{y\in\mathcal{Y}_{0}}p^{(y)}_{i}$.
  2. 2.
    Non trivial classification: $q_{i}=\min_{y\in[K]}p^{(y)}_{i}$; or at most $k_{0}$-sized classification: $q_{i}$ is the $(K-k_{0})$-th smallest element in the set $\{p^{(y)}_{i},y\in[K]\}$.
@@
Let two subset collections $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ that satisfy Assumption 2 with adjusted $p$-values given by $(q_{1,i})_{i\in[m]}$ and $(q_{2,i})_{i\in[m]}$, respectively. Then we can easily check that the intersected collection
:=$\mathcal{I}:=\{I_{1}\cap I_{2},I_{1}\in\mathcal{I}_{1},I_{2}\in\mathcal{I}_{2}\}$
also satisfies Assumption 2 with adjusted $p$-values given by $q_{i}=\max\{q_{1,i},q_{2,i}\}$, $i\in[m]$. This is useful to combine the constraints imposed by the informativeness. For instance, in the regression case (Examples 1.2 and 2.2), we can declare a subset as informative if it excludes a null class while it is of cardinality at most $1$ (see § 6 for a concrete application).
@@
In cases where being informative is linked to a particular location in $\mathcal{Y}$, this testing problem takes an especially meaningful form:
  • Excluding one class in classification: $q_{i}=p^{(y_{0})}_{i}$ tests $H_{0,i}$: “$Y_{n+i}=y_{0}$” versus $H_{1,i}:$$Y_{n+i}\neq y_{0}$”;
  • Excluding $\mathcal{Y}_{0}=[a,b]$ in regression: $q_{i}$ tests $H_{0,i}$: “$Y_{n+i}\in[a,b]$” versus $H_{1,i}:$$Y_{n+i}\notin[a,b]$”.
Note that in case where informative sets are those with small size (e.g., non-trivial classification or length-restricted intervals), we have $\cup_{C\in\mathcal{I}}C=\mathcal{Y}$, so the null hypothesis is always false, and the testing problem is not interesting.
@@
Assumption 2 implies that the $\mathcal{I}$-adjusted $p$-value vector $\mathbf{q}$ is a nondecreasing function of the $p$-value collection, see Lemma A.5; Assumption 2 (iii) is always satisfied up to taking as prediction sets $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ the convex hull of the set in the right-hand-side of (4). It is also satisfied without modifying $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ for any score function such that all the sets $\{y\in\mathbb{R}\>:\>S_{y}(x)\leq s\}$ are intervals, which is often the case, see Remark 2.1 and § 4.
@@ 2.3 Aim: informative selection with FCR guarantees
for the FCP in (2) The FCR expression in (16) for quantifying the errors among the selected is classical in the selective inference literature (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013), since $(Y_{n+i})_{i\in[m]}$ are fixed so this is the false coverage rate on the parameters. On the other hand, in (15), we consider the false coverage rate of prediction intervals. So the false coverage rate is on random outcomes (i.e., $(Y_{n+i})_{i\in[m]}$ in our setting) rather than on parameters, which is related to the Bayes FDR criterion, see, e.g., Efron et al., (2001).
@@
Obviously, the control (18) (class-conditional control, only considered for classification) is stronger than (17) (unconditional control, considered both for classification and regression).
To balance with FCR control, we also consider a notion of power that is the “resolution adjusted” correct decision number:
@@
Hence, for the same selection set, a decision with a smaller covering decision set $\mathcal{C}_{n+i}$ yields higher power. Overall, we aim to maximize the power while controlling the FCR at a pre-specified level $\alpha$. This corresponds to the situation where the analyst would like to select as many examples as possible that are informative, with as narrow as possible a prediction set for the selected examples, while controlling the FCR.
@@ 3.1 Informative selective prediction sets (InfoSP)
In order to have a level $\alpha$ FCR guarantee on the selected examples, it is necessary to correct the threshold $\alpha$ in $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ (4). A standard approach in the selective inference literature (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013) is to use the reduced level $\alpha|\mathcal{S}|/m$ for a selection set $\mathcal{S}$. In order for the selective prediction set procedure to be $\mathcal{I}$ informative, the selection rule needs to be carefully selected. We shall use the following basic observation: selection by a thresholding rule on the family $\mathbf{q}=(q_{i},i\in[m])$ given by (11) will result in selected examples that are $\mathcal{I}$ informative, since $\mathcal{C}^{t}_{n+i}\in\mathcal{I}$ if and only if $q_{i}\leq t$ for all $t\in(0,1]$ by the definition of $q_{i}$. Combining the standard approach for FCR control with this basic observations, leads us to use the following selection thresholding rule which is necessarily $\mathcal{I}$-informative: all examples with $q_{i}$ at most
@@
The informative selective prediction set procedure (InfoSP) based on a $p$-value family $\mathbf{p}=(p^{(y)}_{i},y\in\mathcal{Y},i\in[m]\}$, is defined as $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})=(\mathcal{C}^{\alpha|\mathrm{BH}(\mathbf{q})|/m}_{n+i}(\mathbf{p}))_{i\in\mathrm{BH}(\mathbf{q})}$, that is, is given as follows:
@@
The key theoretical difficulty in proving that the FCR of InfoSP is at most $\alpha$, is that the conformal $p$-values are dependent, and that the selection step is also $p$-value-based (by contrast with Bao et al., (2024)). Thus, the error rate in (17) for the iid model and (18) for the class-conditional model may not be controlled. Bates et al., (2021); Marandon et al., (2024) showed that, for outlier detection, the special positive dependency between the conformal $p$-values is such that the BH procedure remains valid for FDR control. Their results are on a different set of conformal $p$-values, but we develop a similar result for our problem, which enables us to establish the desired error guarantee for various selection rules.
@@
Consider score functions satisfying Assumption 1, an informative subset collection $\mathcal{I}$ satisfying Assumption 2 and a $p$-value collection $\mathbf{p}$ being either $\mathbf{\overline{p}}$ or $\mathbf{{\widetilde{p}}}$, then the information selective prediction set procedure $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})$ defined in Definition 3 controls the FCR at level $\alpha$, respectively in the iid model ($\mathbf{p}=\mathbf{\overline{p}}$) with the control (17) or the class-conditional model ($\mathbf{p}=\mathbf{{\widetilde{p}}}$) with the control (18).
A proof is provided in § C.1, which relies on a more general result, see Theorem A.1 in § A. Theorem A.1 extends the FCR guarantee to more general $p$-value collections (Assumption 4) and more general selection rules (Assumption 5). In particular, it is useful to note that our result applies beyond informative selection (e.g., selection by $p$-value thresholding) and for the two $p$-value collections $\mathbf{\overline{p}}$ and $\mathbf{{\widetilde{p}}}$ under less restrictive conditions that those considered in Theorem 3.1: the independence assumption can be relaxed to an exchangeable assumption (Propositions A.2 and A.3), while the score function can take a general form that can provide an extra improvement (Assumption 6).
Applying Lemma 2.1, we obtain the following theoretical guarantee on the selection rule of InfoSP.
@@
In the setting of Theorem 3.1, the attached selection rule $\mathrm{BH}(\mathbf{q})$ of $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})$ controls the FDR given either by (21) or (22) at level $\alpha$, respectively in the iid model ($\mathbf{p}=\mathbf{\overline{p}}$) or the class-conditional model ($\mathbf{p}=\mathbf{{\widetilde{p}}}$).
@@
For InfoSP, we can avoid the computation of $\mathbf{q}$ by using that the BH procedure is the iterative limit of a recursion (Gao et al.,, 2023). Indeed, since $\mathcal{C}^{\alpha k/m}_{n+i}\in\mathcal{I}$ if and only if $q_{i}\leq\alpha k/m$, the selection $S=\mathrm{BH}(\mathbf{q})$ can be obtained as follows:
  • Step 1: $\mathcal{S}_{1}=\{i\in[m]\>:\>\mathcal{C}^{\alpha}_{n+i}\in\mathcal{I}\}$;
  • Step $t\geq 1$: $\mathcal{S}_{t}=\{i\in\mathcal{S}_{t-1}\>:\>\mathcal{C}^{\alpha|\mathcal{S}_{t-1}|/m}_{n+i}\in\mathcal{I}\}$;
  • Consider $t_{0}$ the first $t$ where $\mathcal{S}_{t}=\mathcal{S}_{t-1}$ and let $\mathcal{S}=\mathcal{S}_{t_{0}}$.

### v3

@@ 2 Preliminaries
2.1 Notation
Expectations and probabilities are denoted for the iid model with $\mathbb{E}_{(X,Y)\sim P_{XY}}(\cdot)$ and $\mathbb{P}_{(X,Y)\sim P_{XY}}(\cdot)$, and for the class-conditional model with $\mathbb{E}_{X\sim P_{X|Y}}(\cdot)$ and $\mathbb{P}_{X\sim P_{X|Y}}(\cdot)$, respectively. If the data generation process is clear from the context, or if the expression is relevant for both models, then the subscript is omitted. In addition, $A\subset B$ means that the set $A$ is included in the set $B$ with a possible equality. For any subset $C\subset\mathbb{R}$, $|C|$ denotes the cardinality of $C$ if the set $C$ is finite, and the Lebesgue measure of $C$ if $C$ is an interval. Finally, for two samples $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, $\mathcal{D}_{1}\cup\mathcal{D}_{2}$ denotes the new sample formed by concatenating the elements of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
@@ 2.2 Conformal prediction sets
where $\mathbf{p}=(p^{(y)}_{i},y\in\mathcal{Y},i\in[m])$ is a collection of conformal $p$-values satisfying the following super-uniform guarantee:
@@
where the above probability is computed either in the iid ($Y_{n+i}$ random) or conditional ($Y_{n+i}$ fixed) model. Super-uniformity (5) implies directly that $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ in (4) provides $1-\alpha$ coverage for $Y_{n+i}$, that is,
@@
The $p$-value family is built from the calibration and test sample by using non-conformity score functions $S_{y}:x\in\mathbb{R}^{d}\mapsto\mathbb{R}$, $y\in\mathcal{Y}$, measuring the inadequacy between $y$ and the prediction at point $x$. Importantly, we follow a split/inductive conformal approach, where the score functions have been computed from an independent training data sample so that they can be considered as fixed here [^3] Note that we can relax slightly this assumption: our theory also allows this function to depend on the calibration plus test samples in an exchangeable way, see Assumption 6. (and all probabilities/expectations are taken conditional on that training sample).
@@
The score functions $S_{y}(\cdot)$, $y\in\mathcal{Y}$, have been computed from an independent training sample and the computed scores $(S_{Y_{i}}(X_{i}),i\in[n+m])$ have no ties almost surely. When $\mathcal{Y}=\mathbb{R}$, the score function is regular in the following sense: for every $x\in\mathbb{R}^{d}$, the function $y\in\mathbb{R}\mapsto S_{y}(x)\in\mathbb{R}$ is right continuous with left limits.
@@
The prediction set $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ in (4) can be described as a score level set, with a threshold depending on the calibration scores. Formally, we have
@@
where the score threshold $\hat{s}_{\alpha}$ is $S_{(\lceil(1-\alpha)(n_{\tiny\mbox{cal}}+1)\rceil)}$ with $S_{(1)}\leq\dots\leq S_{(n_{\tiny\mbox{cal}})}$ being the ordered calibration scores (and with the convention $S_{(n_{\tiny\mbox{cal}}+1)}=+\infty$). For full-calibrated $p$-values, the $n_{\tiny\mbox{cal}}=n$ calibration scores are $S_{Y_{j}}(X_{j})$, $j\in[n]$. For class-conditional $p$-values, the $n_{\tiny\mbox{cal}}=|\mathcal{D}^{(y)}_{{\tiny\mbox{cal}}}|$ calibration scores are $S_{y}(X_{j})$, $j\in\mathcal{D}^{(y)}_{{\tiny\mbox{cal}}}$ ($\hat{s}_{\alpha}$ depends on $y$).
@@ 2.3 $\mathcal{I}$-adjusted $p$-values
    If a prediction set is informative, then all the prediction sets it contains are also informative, that is, for all $\mathcal{C},\mathcal{C}^{\prime}$ (subsets of $[K]$ for classification, intervals of $\mathbb{R}$ for regression) with $\mathcal{C}^{\prime}\subset\mathcal{C}$, $\mathcal{C}\in\mathcal{I}$ implies $\mathcal{C}^{\prime}\in\mathcal{I}$.
@@
    (for regression) For all $\alpha\in(0,1)$, almost surely, $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ is an interval of $\mathbb{R}$.
@@
with by convention $q_{i}=1$ if the set is empty. Assumption 2 can be easily checked for Examples 1.1 and 1.2, with an explicit expression for $q_{i}$’s (see Section F for the detailed derivations of these expressions).
@@
For regression (see § 4 for more details): $q_{i}=\sup_{y\in[a,b]}p^{(y)}_{i}$ for intervals excluding $\mathcal{Y}_{0}=[a,b]$; $q_{i}=(n+1)^{-1}\big{(}1+\sum_{j=1}^{n}{\mathbf{1}\{S_{Y_{j}}(X_{j})>A\}}\big{)}$ for length-restricted intervals $I$ (with $|I|\leq 2\lambda_{0}$), where $A=\lambda_{0}/\sigma(X_{n+i})$ for $S_{y}(x)=|y-\mu(x)|/\sigma(x)$ and $A=\lambda_{0}-(q_{\beta_{1}}(X_{n+i})-q_{\beta_{0}}(X_{n+i}))/2$ for $S_{y}(x)=\max(q_{\beta_{0}}(x)-y,y-q_{\beta_{1}}(x))$.
@@
For classification: $q_{i}=p^{(y_{0})}_{i}$ for excluding one-class; $q_{i}=\max_{y\in\mathcal{Y}_{0}}p^{(y)}_{i}$ for excluding several classes; $q_{i}=\min_{y\in[K]}p^{(y)}_{i}$ for non-trivial classification; $q_{i}=$ the $(K-k_{0})$-th smallest element in the set $\{p^{(y)}_{i},y\in[K]\}$ for at most $k_{0}$-sized classification.
@@
Let $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ be two subset collections that satisfy Assumption 2 with adjusted $p$-values given by $(q_{1,i})_{i\in[m]}$ and $(q_{2,i})_{i\in[m]}$, respectively. Then we can easily check that the intersected collection :=$\mathcal{I}:=\{I_{1}\cap I_{2},I_{1}\in\mathcal{I}_{1},I_{2}\in\mathcal{I}_{2}\}$ also satisfies Assumption 2 with adjusted $p$-values given by $q_{i}=\max\{q_{1,i},q_{2,i}\}$, $i\in[m]$. This is useful to combine the constraints imposed by the informativeness. For instance, in the classification case (Examples 1.2 and 2.2), we can declare a subset as informative if it excludes a null class while it is of cardinality at most one (see § B for a concrete application).
@@
More specifically, using the $p$-value $q_{i}$, $H_{0,i}$ is rejected at level $\alpha$ whenever $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})\in\mathcal{I}$, and it satisfies a valid super-uniformity property, see Remark 2.4.
In cases where being informative is linked to particular values in $\mathcal{Y}$, this testing problem takes an especially meaningful form. In classification, for excluding one class in classification: $q_{i}=p^{(y_{0})}_{i}$ tests $H_{0,i}$: “$Y_{n+i}=y_{0}$” versus $H_{1,i}:$$Y_{n+i}\neq y_{0}$”. In regression, for excluding $\mathcal{Y}_{0}=[a,b]$: $q_{i}$ tests $H_{0,i}$: “$Y_{n+i}\in[a,b]$” versus $H_{1,i}:$$Y_{n+i}\notin[a,b]$”. Note that in case where informative sets are those with small size (e.g., non-trivial classification or length-restricted intervals), we have $\cup_{C\in\mathcal{I}}C=\mathcal{Y}$, so the null hypothesis is always false, and the testing problem (12) is not interesting.
@@
Assumption 2 implies that the $\mathcal{I}$-adjusted $p$-value vector $\mathbf{q}$ is a nondecreasing function of the $p$-value collection, see Lemma C.5; Assumption 2 (iii) is always satisfied up to taking as prediction sets $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ the convex hull of the set in the right-hand-side of (4). It is also satisfied without modifying $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ for any score function such that all the sets $\{y\in\mathbb{R}\>:\>S_{y}(x)\leq s\}$ are intervals, which is often the case, see Remark 2.1 and § 4.
Remark 2.3.
In a context of building online confidence intervals, Weinstein and Ramdas, (2020) have proposed to report only intervals that are “good” in the sense that they “localize” the signal. Formalizing their proposal in our regression setting and with our notation, this corresponds to consider the informative collection =$\mathcal{I}=\{\mbox{$I$ interval of $\mathbb{R}$}\>:\>I\subset\mathcal{C}_{\ell}\mbox{ for one $\ell\in[L]$}\},$ where $\mathcal{C}_{\ell}$, $\ell\in[L]$, are pre-specified disjoint subsets of $\mathbb{R}$. This collection satisfies Assumption 2 and the corresponding $\mathcal{I}$-adjusted $p$-values are given by $q_{i}=\min_{\ell\in[L]}\sup_{y\notin\mathcal{C}_{\ell}}p_{i}^{(y)}$.
Remark 2.4.
The following super-uniformity property holds for the $\mathcal{I}$-adjusted $p$-value $q_{i}$ (11) and the null $H_{0,i}$ (12):
$P(q_{i}\leq\alpha,\mbox{$H_{0,i}$ true})\leq\alpha,\>\>\mbox{ for all $\alpha\in(0,1)$.}$
To see this, note that when $H_{0,i}$ is true, we have $Y_{n+i}\in\cap_{C\in\mathcal{I}}C^{c}$. From the definition, $q_{i}\leq\alpha$ means $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})\in\mathcal{I}$. Therefore, if both $q_{i}\leq\alpha$ and $H_{0,i}$ is true, we have $Y_{n+i}\notin\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$, i.e., that $p^{(Y_{n+i})}_{i}\leq\alpha$, which occurs with probability at most $\alpha$ by (5).
@@ 2.4 Aim: informative selection with FCR guarantees
for the FCP in (2). The FCR expression in (14) for quantifying the errors among the selected is classical in the selective inference literature: since $(Y_{n+i})_{i\in[m]}$ is fixed, this is the false coverage rate on the parameters (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013). On the other hand, in (13), the false coverage rate is on random outcomes (i.e., $(Y_{n+i})_{i\in[m]}$ in our setting) rather than on parameters, which is the usual setting in selective conformal inference (see references in § 1) and is related to the Bayes FDR criterion in the multiple testing literature, see, e.g., Efron et al., (2001).
@@
Obviously, the class-conditional control (16) (considered only for classification) is stronger than the unconditional control (15) (considered both for classification and regression).
To balance with FCR control, we also consider the resolution adjusted power:
@@
Hence, for the same selection set, a decision with a smaller covering decision set $\mathcal{C}_{n+i}$ yields higher power. Our aim is to maximize the resolution adjusted power (i.e., informally, to select as many examples as possible that are informative, with as narrow as possible a prediction set for the selected examples), while controlling the FCR at a pre-specified level $\alpha$.
@@ 3.1 Informative selective prediction sets (InfoSP)
In order to have a level $\alpha$ FCR guarantee on the selected examples, it is necessary to correct the threshold $\alpha$ in $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ (4). A standard approach in the selective inference literature (Benjamini and Yekutieli,, 2005; Benjamini and Bogomolov,, 2013) is to use the reduced level $\alpha|\mathcal{S}|/m$ for a selection set $\mathcal{S}$. In order for the selective prediction set procedure to be $\mathcal{I}$ informative, the selection rule needs to be carefully selected. We shall use the following basic observation: selection by a thresholding rule on the family $\mathbf{q}=(q_{i},i\in[m])$ given by (11) will result in selected examples that are $\mathcal{I}$ informative for prediction sets that are constructed at a level that is at least at the selection threshold, since $\mathcal{C}^{t}_{n+i}\in\mathcal{I}$ if and only if $q_{i}\leq t$ for all $t\in(0,1]$ by the definition of $q_{i}$. Combining the standard approach for FCR control with this basic observation, leads us to use the following selection thresholding rule which is necessarily $\mathcal{I}$-informative: all examples with $q_{i}$ at most
@@
The informative selective prediction set procedure (InfoSP) based on a $p$-value family $\mathbf{p}=(p^{(y)}_{i},y\in\mathcal{Y},i\in[m])$, is defined as $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})=(\mathcal{C}^{\alpha|\mathrm{BH}(\mathbf{q})|/m}_{n+i}(\mathbf{p}))_{i\in\mathrm{BH}(\mathbf{q})}$, that is, is given as follows:
@@
The key theoretical difficulty in proving that the FCR of InfoSP is at most $\alpha$, is that the conformal $p$-values are dependent, and that the selection step is also $p$-value-based (by contrast with Bao et al.,, 2024). Thus, the error rate in (13) for the iid model and (14) for the class-conditional model may not be controlled. Bates et al., (2023); Marandon et al., (2024) showed that, for outlier detection, the special positive dependency between the conformal $p$-values is such that the BH procedure remains valid for FDR control. Their results are on a different set of conformal $p$-values, but we develop a similar result for our problem, which enables us to establish the desired error guarantee for various selection rules.
@@
Consider score functions satisfying Assumption 1, an informative subset collection $\mathcal{I}$ satisfying Assumption 2 and a $p$-value collection $\mathbf{p}$ being either $\mathbf{\overline{p}}$ (full-calibrated) or $\mathbf{{\widetilde{p}}}$ (class-calibrated), then the $\mathcal{I}$-informative selective prediction set procedure $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})$ (Definition 1) controls the FCR at level $\alpha$, respectively in the iid model ($\mathbf{p}=\mathbf{\overline{p}}$) with the control (15) or the class-conditional model ($\mathbf{p}=\mathbf{{\widetilde{p}}}$) with the control (16).
A proof is provided in § E.1, which relies on a more general result, see Theorem C.1. The latter provides FCR guarantee to more general $p$-value collections (see Assumption 4). For instance, it is valid for the two $p$-value collections $\mathbf{\overline{p}}$ and $\mathbf{{\widetilde{p}}}$ under less restrictive conditions that those considered in Theorem 3.1: the independence assumption can be relaxed to an exchangeable assumption (Propositions C.2 and C.3), while the score function can take a general form that can provide an extra improvement (Assumption 6). In addition, our result applies beyond informative selection, to any concordant selection rule (see Assumption 5), as defined in Benjamini and Yekutieli, (2005); Benjamini and Bogomolov, (2013). For instance, it applies for any monotone $p$-value-based thresholding rule (see Section C.4 for more details). This is in sharp contrast with the work of Bao et al., (2024) that does not provide finite sample FCR control at the desired level for $p$-value-based selection rules, because the selection rule therein should be independent of the calibration scores.
Let us also mention that an alternative proof of Theorem 3.1 can be obtained by using the property that a particular $p$-value based family is PRDS (positive regression dependence on each one from a subset Benjamini and Yekutieli,, 2001), see Section G. This can be seen as an extension of Lemma 5 (iii) in Jin and Candes, (2023) to our general informative setting.
We apply Lemma 2.1 to obtain the following FDR guarantee for InfoSP.
@@
In the setting of Theorem 3.1, the selection rule $\mathrm{BH}(\mathbf{q})$ of $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p})$ provides level $\alpha$ FDR control, given either by (19) in the iid model ($\mathbf{p}=\mathbf{\overline{p}}$) or (20) in the class-conditional model ($\mathbf{p}=\mathbf{{\widetilde{p}}}$).
@@
For InfoSP, we can avoid the computation of $\mathbf{q}$ by using that the BH procedure is the iterative limit of a recursion (Gao et al.,, 2023). Indeed, since $\mathcal{C}^{\alpha k/m}_{n+i}\in\mathcal{I}$ if and only if $q_{i}\leq\alpha k/m$, the selection $S=\mathrm{BH}(\mathbf{q})$ can be obtained as follows: Step 1: $\mathcal{S}_{1}=\{i\in[m]\>:\>\mathcal{C}^{\alpha}_{n+i}\in\mathcal{I}\}$; Step $t\geq 1$: $\mathcal{S}_{t}=\{i\in\mathcal{S}_{t-1}\>:\>\mathcal{C}^{\alpha|\mathcal{S}_{t-1}|/m}_{n+i}\in\mathcal{I}\}$; Consider $t_{0}$ the first $t$ where $\mathcal{S}_{t}=\mathcal{S}_{t-1}$ and let $\mathcal{S}=\mathcal{S}_{t_{0}}$.